The pH of a 0.30 M solution of a weak base is 10.82. What is the Kb of the base?
Answer #1

pOH =14- pH= 14-10.82=3.81

pOH = - log[OH-]

= -log 10-3.81

= - log 1.55 ×10-4

[OH-] = 1.55× 10^-4

[OH-] =√KbC,

C= .3 M

Kb=[OH-]^2/C=( 1.55×10^-4)^2/.3

= 8 × 10^-8